OpenStudy (anonymous):

represents the curve given by x = t ^ 2 - 2t, y = t find the area under the curve from t = 2 to t = 3

OpenStudy (anonymous):

Hi Kevindorado Do you know how to find the Area of a curve using Integration? An idea will be to diffentiate x wrt to t. Since the limits are expressed in t, its best to integrate with t. Are you able to take it from there?

OpenStudy (anonymous):

Area = \[\int\limits_{2}^{3} y dx\] y=t x=t^2 dx = 2t dt so, Area = \[\int\limits_{2}^{3} (t)(2t) dt\] = \[\int\limits_{2}^{3} 2t ^{2}\] i think you can solve it from here onwards.

OpenStudy (anonymous):

*dt